OpenStudy (anonymous):

Solve for B. A = 3(B + C) B = (A - C)/3 B = (3A - C)/3 B = (A - 3C)/3

OpenStudy (anonymous):

A = 3(B+C) A/3 = B+C B = A/3 - C

OpenStudy (anonymous):

Same answer though